# An Accelerated Hybrid Low-Rank Augmented Lagrangian Method for Large-Scale Semidefinite Programming on GPU

### v1

@@ 4 Numerical Experiments
In this section, we provide provide extensive experiments comparing cuHALLaR, HALLaR, and cuLoRADS. The experiments concern SDP instances for the matrix completion, the maximum stable set, and the phase retrieval problems. These problems were also used in the benchmark of the original HALLaR paper [47].
@@ 4.1 Experimental Setup
We implemented cuHALLaR as a Julia [3] module. Our implementation uses the CUDA language [13] for interfacing with NVIDIA CUDA GPUs. The experiments are performed running CUDA 12.6.1 and Julia 1.11.3. The reported running times do not take into account the pre-compilation time.
@@
Given a tolerance $\epsilon>0$, all methods stop when an iterate $(X,p)\in\Delta^{n}\times\mathbb{R}^{m}$ satisfying
$\max\left\{\frac{\|\mathcal{A}X-b\|_{2}}{1+\|b\|_{1}},\frac{|\mathrm{pval}-\mathrm{dval}|}{1+|\mathrm{pval}|+|\mathrm{dval}|},\frac{\max\{0,-\lambda_{\min}(C+{\mathcal{A}}^{*}{p})\}}{1+\|C\|_{1}}\right\}\leq\epsilon.$ (15)
is obtained, where $\mathrm{pval}=C\bullet X$ and $\mathrm{dval}=-b^{\top}p$ denote the primal and dual values of $(X,p)$, respectively.
Scaling.
We have assumed in Section 1 that the trace of $X$ is bounded by one. However, cuHALLaR and HALLaR allow $\operatorname*{tr}X\leq 1$ to be replaced by the more general constraint $\operatorname*{tr}(X)\leq\tau$ for some $\tau>0$. This is done by reformulating the SDP as in (P) using the changes of variables $X\leftarrow X/\tau$ and changing the rhs as $b\leftarrow b/\tau$. (15) is used in the experiments of phase retrieval and matrix completion.
@@
HALLaR and cuHALLaR are initialized using an initial random $n\times 1$ matrix with entries generated independently from the Gaussian distribution over $\mathbb{R}$ (or over $\mathbb{C}$ for complex-valued SDPs) which is then scaled to be a unit vector. The dual initial point $p_{0}$ is taken to be the zero vector.
@@
HALLaR and cuHALLaR require the following user inputs to describe the SDP instance:
  1. (i)
    The vector $b\in\mathbb{R}^{m}$ and the scalar $\tau>0$. Here, $\tau$ represents the upper bound for the trace of the primal matrix variable $X\in\mathbb{S}^{n}_{+}$, such that $\mathrm{Tr}(X)\leq\tau$.
  2. (ii)
    A routine for evaluating $CU$ for arbitrary matrices $U$ of compatible dimensions.
  3. (iii)
    A routine for evaluating $(\mathcal{A}^{*}p)U$ for arbitrary vectors $p$ and matrices $U$ of compatible dimensions.
  4. (iv)
    A routine for evaluating $\mathcal{A}(UU^{\top})$ for arbitrary matrices $U$ of compatible dimensions.
We are aware that the above operator-based input format for SDPs is not prevalent as matrix-based representations such as the SDPA format [72, 73] or the SEDUMI format [59]. Matrix-based representations have been widely adopted by numerous SDP solvers, and are quite effective for SDPs involving sparse matrices. However, these representations might not be effective for SDPs involving dense matrices. In particular, the stable set and the phase retrieval SDPs involve fully dense matrices, and any code that requires input in a matrix-based format (e.g., SDPA format) can take a long time to read and manipulate their data. In contrast, our operator-based input format enables us to efficiently handle extremely large stable set and phase retrieval SDP instances.
Ratio:
In all tables below, the entry labeled “Ratio” reports the running time of the CPU implementation divided by the running time of the GPU implementation (or, when comparing two solvers, the time of the first solver over the time of the second). For instance, in Table 2 the first row shows cuHALLaR runtime of 0.77 sec and a cuLoRADS runtime of 123.71 sec, yielding a ratio of 160.662. This means that cuLoRADS took 160.662 times longer to solve that instance than cuHALLaR.
@@ 4.2 Experiments for matrix completion
$\min_{X\in\mathbb{S}^{n_{1}+n_{2}}}\Bigl{\{}\,\frac{1}{2}\mathrm{tr}(X):X=\begin{pmatrix}W_{1}&Y\\ Y^{\top}&W_{2}\end{pmatrix}\succeq 0\,,\quad Y_{i,j}=M_{i,j}\,,\;\forall\,(i,j)\in\Omega).\Bigr{\}}$ (16)
@@
Matrix completion instances are generated randomly, using the following procedure: Given $r\leq n_{1}\leq n_{2}$, the hidden solution matrix $M$ is the product $UV^{\top}$, where the matrices $U\in\mathbb{R}^{n_{1}\times r}$ and $V\in\mathbb{R}^{n_{2}\times r}$ have independent standard Gaussian random variables as entries. Afterward, $m$ independent and uniformly random entries from $M$ are taken, where $m=\lceil\gamma\ r(n_{1}+n_{2}-r)\rceil$ and $\gamma=r\log(n_{1}+n_{2})$ is the oversampling ratio.
Problem Instance Runtime (seconds) / rank
Size $(n_{1},n_{2};m)$ $r$ cuHALLaR cuLoRADS Ratio
3,000, 7,000; 828,931 3 0.77/3 123.71/19 160.662
3,000, 7,000; 828,931 3 0.7/3 11.42/19 16.314
3,000, 7,000; 2,302,586 5 1.28/5 3.19/19 2.492
3,000, 7,000; 2,302,586 5 1.09/5 2.26/19 2.073
10,000, 21,623; 2,948,996 3 1.22/3 5.14/21 4.213
10,000, 21,623; 8,191,654 5 1.95/5 4.71/21 2.415
25,000, 50,000; 3,367,574 2 1.2/2 10.77/23 8.975
25,000, 50,000; 7,577,040 3 1.82/3 21.31/23 11.708
30,000, 70,000; 4,605,171 2 1.63/2 17.02/24 10.441
30,000, 70,000; 10,361,633 3 2.57/3 18.45/24 7.178
50,000, 100,000; 7,151,035 2 1.89/2 23.67/24 12.523
50,000, 100,000; 16,089,828 3 3.68/3 27.95/24 7.595
80,000, 120,000; 9,764,859 2 2.45/2 20.66/25 8.432
80,000, 120,000; 21,970,931 3 5.02/3 32.28/25 6.430
80,000, 120,000; 7,996,791 3 2.80/3 17.94/25 6.407
120,000, 180,000; 11,996,886 3 3.74/3 36.47/25 9.751
160,000, 240,000; 15,996,836 3 4.90/3 52.61/25 10.736
240,000, 360,000; 23,996,846 3 7.03/3 86.80/25 12.347
Table 2: cuHALLaR vs cuLoRADS runtimes for the matrix completion problem. Seed was 0 for all instances. Tolerance is set at $10^{-5}$.
Table 2 provides a direct comparison between cuHALLaR and cuLoRADS on matrix completion problems, revealing cuHALLaR’s substantial performance advantages. Our GPU implementation consistently outperforms cuLoRADS across all problem instances with speedups ranging from 2x to over 160x. Most notably, cuHALLaR solves these problems at a significantly lower rank, demonstrating its superior ability to find optimal low-rank solutions efficiently. This comparison highlights an advantage of our approach: cuLoRADS struggles with larger problems due to its reliance on the SDPA format and less efficient memory handling, while cuHALLaR’s operator-based formulation enables it to tackle much larger instances without encountering the same memory limitations.
Problem Instance Runtime (seconds) / rank
Size $(n_{1},n_{2};m)$ $r$ HALLaR cuHALLaR Ratio
3,000, 7,000; 828,931 3 7.63/3 0.77/3 9.91
3,000, 7,000; 828,931 3 7.52/3 0.70/3 10.74
3,000, 7,000; 2,302,586 5 34.52/5 1.28/5 26.97
3,000, 7,000; 2,302,586 5 34.42/5 1.09/5 31.58
10,000, 21,623; 2,948,996 3 38.88/3 1.22/3 31.87
10,000, 21,623; 2,948,996 3 34.80/3 1.11/3 31.35
10,000, 21,623; 8,191,654 5 141.17/5 2.05/5 68.86
10,000, 21,623; 8,191,654 5 140.55/5 1.95/5 72.08
25,000, 50,000; 3,367,574 2 43.52/2 1.20/2 36.27
25,000, 50,000; 7,577,040 3 79.93/3 1.82/3 43.92
30,000, 70,000; 4,605,171 2 60.57/2 1.63/2 37.16
30,000, 70,000; 10,361,633 3 145.76/3 2.57/3 56.72
50,000, 100,000; 7,151,035 2 121.63/2 1.89/2 64.35
50,000, 100,000; 16,089,828 3 242.00/3 3.68/3 65.76
80,000, 120,000; 9,764,859 2 237.38/2 2.45/2 96.89
80,000, 120,000; 21,970,931 3 490.07/3 5.02/3 97.62
80,000, 120,000; 7,996,791 3 217.97/3 2.80/3 77.85
120,000, 180,000; 11,996,886 3 379.99/3 3.74/3 101.60
160,000, 240,000; 15,996,836 3 430.87/3 4.90/3 87.93
240,000, 360,000; 23,996,846 3 834.17/3 7.03/3 118.66
320,000, 480,000; 31,996,932 3 1,152.68/3 9.60/3 120.07
400,000, 600,000; 39,996,783 3 1,587.08/3 13.05/3 121.62
480,000, 720,000; 47,996,814 3 2,299.54/4 16.06/3 143.18
560,000, 840,000; 55,996,812 3 2,126.11/4 18.23/3 116.63
640,000, 960,000; 63,996,921 3 2,659.69/3 20.53/3 129.55
720,000, 1,080,000; 71,996,873 3 2,904.02/3 26.90/3 107.96
800,000, 1,200,000; 79,996,838 3 2,833.93/3 27.05/3 104.77
1,600,000, 2,400,000; 159,996,869 3 8,364.62/3 60.68/3 137.85
2,400,000, 3,600,000; 239,996,782 3 Timed Out 100.17/3 N/A
3,200,000, 4,800,000; 319,996,871 3 Timed Out 142.27/3 N/A
@@
Table 3: Performance comparison between HALLaR (CPU) and cuHALLaR for matrix completion problems. A relative tolerance of $\epsilon=10^{-5}$ is used throughout. Ratio is CPU Time / GPU Time. Time limit was set at 12 hours.
Table 3 highlights the speedup achieved by cuHALLaR over its CPU counterpart. The GPU implementation consistently delivers 10-140x faster solution times across standard problem sizes while maintaining identical solution quality. Below the midrule, we showcase truly large-scale instances with dimensions from $(n_{1},n_{2})=(320,000,480,000)$ up to $(3,200,000,4,800,000)$. For these instances, cuHALLaR achieves the most remarkable speedups of 104-143x. Most significantly, for problems with dimensions $(n_{1},n_{2})=(2,400,000,3,600,000)$ and higher, the CPU implementation times out after 12 hours while cuHALLaR solves them in under 2.5 minutes. These “extra-large” instances couldn’t be evaluated on cuLoRADS due to its reliance on the SDPA format, which becomes prohibitively memory-intensive for dense problems of this scale. This demonstrates not only the efficiency of cuHALLaR, but also its ability to handle problem sizes previously deemed computationally intractable and would have taken hours on the original CPU HALLaR implementation.
Note that the experimental results in Table 3 are divided into two distinct sets of problem instances. The first set showcases standard problem dimensions where both CPU and GPU implementations can successfully operate, allowing for direct performance comparison. Below the midrule, we present a set of significantly larger-scale problems with dimensions ranging from $(n_{1},n_{2})=(320,000,480,000)$ up to $(3,200,000,4,800,000)$.
@@ 4.3 Experiments for maximum stable set
Given a graph $G=([n],E)$, the maximum stable set problem consists of finding a subset of vertices of largest cardinality such that no two vertices are connected by an edge. Lovász introduced a constant, the $\vartheta$-function, which upper bounds the value of the maximum stable set. The $\vartheta$-function is the value of the SDP
$\max\quad\{{e}{e}^{\top}\bullet X\;:\;X_{ij}=0,\;ij\in E,\;\mathrm{Tr}(X)=1,\;X\succeq 0,\quad X\in\mathbb{S}^{n}\}$ (17)
where $e=(1,1,\dots,1)\in\mathbb{R}^{n}$ is the all ones vector. It was shown in [28] that the $\vartheta$-function agrees exactly with the stable set number for perfect graphs.
@@
10,937; 75,488 wing_nodal DIMACS10 3,826.34/130 246.71/127 15.51
16,384; 49,122 delaunay_n14 DIMACS10 643.90/36 102.43/34 6.29
16,386; 49,152 fe-sphere DIMACS10 21.74/3 17.82/3 1.22
22,499; 43,858 cs4 DIMACS10 749.95/15 588.46/115 1.27
25,016; 62,063 hi2010 DIMACS10 2,607.24/15 132.85/21 19.63
@@
32,580; 77,799 vt2010 DIMACS10 2,952.84/18 196.59/60 15.02
48,837; 117,275 nh2010 DIMACS10 7,694.06/20 2,186.74/172 3.52
24,300; 34,992 aug3d GHS_indef 30.10/1 2.08/1 14.47
32,430; 54,397 ia-email-EU Network Repo 598.76/5 52.86/4 11.33
11,806; 32,730 Oregon-2 SNAP 1,581.11/23 467.75/55 3.38
21,363; 91,286 ca-CondMat SNAP 8,521.04/59 348.59/77 24.44
31,379; 65,910 as-caida_G_001 SNAP 2,125.41/8 428.41/27 4.96
26,518; 65,369 p2p-Gnutella24 SNAP 312.75/4 55.99/10 5.59
22,687; 54,705 p2p-Gnutella25 SNAP 242.42/4 88.53/13 2.74
36,682; 88,328 p2p-Gnutella30 SNAP 506.26/5 61.84/10 8.19
62,586; 147,892 p2p-Gnutella31 SNAP 1,484.71/5 245.33/29 6.05
49,152; 69,632 cca AG-Monien 63.09/2 12.30/2 5.13
49,152; 73,728 ccc AG-Monien 14.52/2 20.62/2 0.70
49,152; 98,304 bfly AG-Monien 15.72/2 10.35/2 1.52
16,384; 32,765 debr_G_12 AG-Monien 250.40/10 189.57/10 1.32
32,768; 65,533 debr_G_13 AG-Monien 471.80/9 638.54/13 0.74
65,536; 131,069 debr_G_14 AG-Monien 474.22/9 29.51/9 16.07
131,072; 262,141 debr_G_15 AG-Monien 521.88/10 28.07/10 18.59
262,144; 524,285 debr_G_16 AG-Monien 1,333.31/12 84.28/13 15.82
524,288; 1,048,573 debr_G_17 AG-Monien 6,437.05/12 171.81/13 37.47
1,048,576; 2,097,149 debr_G_18 AG-Monien 16,176.30/13 119.50/13 135.37
Table 4: Runtimes (in seconds) for the Maximum stable set problem. A relative tolerance of $\epsilon=10^{-5}$ is set.
The results in Table 4 demonstrate a strong scaling advantage for GPU implementation as problem size increases. For the largest instances, cuHALLaR achieves a remarkable 135x speedup over CPU implementation. Performance patterns vary across graph families: DIMACS10 instances show substantial speedups up to 19.6x, SNAP graphs deliver consistent 3-24x improvements, and AG-Monien graphs exhibit dramatic acceleration of 16-135x for larger instances. Some smaller or sparser graphs show modest or even negative speedups, highlighting that GPU acceleration benefits depend on sufficient computational complexity to amortize parallelization overhead. These findings reconfirm that GPU acceleration becomes increasingly advantageous as problem dimensions grow, enabling cuHALLaR to efficiently solve million-scale instances that would be computationally prohibitive on CPU alone.
Problem Instance HALLaR cuHALLaR cuLoRADS HALLaR/cuHALLaR cuLoRADS/cuHALLaR
@@
Table 5: Runtimes (in seconds) for the Maximum Stable Set problem on Hamming graphs. A relative tolerance of $\epsilon=10^{-5}$ is set.

### v2

@@ 5 Numerical Experiments
In this section, we provide extensive experiments comparing cuHALLaR, HALLaR, and cuLoRADS on three problem classes: matrix completion, maximum stable set, and phase retrieval. These problems were also used in the benchmark of the original HALLaR paper [50]. The pre-compiled binary solver can be found at https://github.com/OPTHALLaR.
We attempted to run CuClarabel [13] on the matrix completion and stable set experiments, but CuClarabel failed even on relatively small SDPA files. Moreover, CuClarabel’s lack of native SDPA support precluded straightforward benchmarking, so we omit comparisons against it.
The comparisons between cuHALLaR and cuLoRADS evaluate cuHALLaR using the HSLR format described in Section 4 against cuLoRADS using SDPA format. This comparison highlights the performance and memory efficiency advantages of our HSLR representation and cuHALLaR’s implementation.
@@ 5.1 Experimental Setup
We implemented cuHALLaR as a Julia [4] module. Our implementation uses the CUDA platform [14] for interfacing with NVIDIA CUDA GPUs. The experiments are performed using CUDA 12.6.1 and Julia 1.11.3. The reported running times do not take into account the pre-compilation time.
@@
Given a tolerance $\epsilon>0$, both solvers stop when a set of relative error metrics falls below $\epsilon$. For cuHALLaR, an iterate $(X,p,\theta)\in\Delta^{n}_{\tau}\times\mathbb{R}^{m}\times\mathbb{R}_{+}$ is deemed optimal if it satisfies
$\max\left\{\frac{\|\mathcal{A}X-b\|_{2}}{1+\|b\|_{1}},\frac{|\mathrm{pval}-\mathrm{dval}|}{1+|\mathrm{pval}|+|\mathrm{dval}|},\frac{\max\{0,-\lambda_{\min}(S)\}}{1+\|C\|_{1}}\right\}\leq\epsilon,$ (31)
where $\mathrm{pval}=C\bullet X$ is the primal value, $S=C+{\mathcal{A}}^{*}{p}+\theta I$ is the dual slack matrix, and the dual value is $\mathrm{dval}=-b^{\top}p-\tau^{2}\theta$. The solver cuLoRADS employs a similar set of metrics; however, as it addresses a formulation without the spectraplex constraint, its dual value is defined as $\mathrm{dval}=-b^{\top}p$.
@@
Both cuHALLaR and HALLaR are initialized by setting the dual variable to $p_{0}=0$ and the primal factor $U_{0}$ to an $n\times 1$ matrix with entries drawn independently from a standard Gaussian distribution.
@@
cuHALLaR binary accepts two different input formats: HSLR and SDPA formats. For SDP data prepared in SDPA format, the user must also supply the trace bound $\tau>0$ in (2) for the constraint $\mathrm{Tr}(X)\leq\tau^{2}$ externally via the command-line interface or a parameter file. On the other hand, HSLR embeds the trace bound $\tau$ directly within the file itself, so that the user does not need to supply it separately. For more details and examples of preparing data in SDPA and HSLR format, the reader should refer to our user manual [1].
Our computational experiments also report results for HALLaR and cuHALLaR versions (not available to the public) where the data is generated in the memory. Specifically, instead of reading the problem data from an input file, these versions use subroutines for computing $\tilde{\mathcal{A}}(UU^{\top})$ and $\tilde{\mathcal{A}}^{*}(\tilde{p})U$ for any $\tilde{p}=(1,p)\in\mathbb{R}^{m+1}$ and $U\in\mathbb{R}^{n\times r}$. Moreover, these two “on-the-fly” versions have the advantage of being able to run the phase retrieval experiments, which are expensive to store in either SDPA or HSLR.
@@ 5.2 Experiments for matrix completion
$\min_{X\in\mathbb{S}^{n_{1}+n_{2}}}\left\{\,\frac{1}{2}\mathrm{Tr}(X):X=\begin{pmatrix}W_{1}&Y\\ Y^{\top}&W_{2}\end{pmatrix}\succeq 0\,,\quad Y_{i,j}=M_{i,j}\,,\;\forall\,(i,j)\in\Omega)\right\}.$ (32)
@@
Matrix completion instances are generated randomly, using the following procedure: given $r\leq n_{1}\leq n_{2}$, the hidden solution matrix $M$ is the product $UV^{\top}$, where the matrices $U\in\mathbb{R}^{n_{1}\times r}$ and $V\in\mathbb{R}^{n_{2}\times r}$ have independent standard Gaussian random variables as entries. Afterward, $m$ independent and uniformly random entries from $M$ are taken, where $m=\lceil\gamma\ r(n_{1}+n_{2}-r)\rceil$, and $\gamma=r\log(n_{1}+n_{2})$ is the oversampling ratio.
Table 2 compares HALLaR, cuHALLaR, and cuLoRADS on matrix completion problems. The second block of columns demonstrates the tremendous speedups our GPU-accelerated solver cuHALLaR achieves over HALLaR using on-the-fly-problem generation. On many instances, cuHALLaR achieves speedups of 10x to 188x compared to HALLaR. For the largest instance considered with $(n_{1},n_{2})=(800,\!000,1,\!200,\!000)$, cuHALLaR solves the problem in 53 seconds, while HALLaR requires nearly 2.5 hours.
The last block of columns compares cuHALLaR, using both HSLR and SDPA formats via our binary, against cuLoRADS, which uses SDPA format. Here, cuHALLaR with HSLR format consistently outperforms cuLoRADS across all instances, achieving speedups ranging from 0.9x to 37x. Notably, cuHALLaR efficiently finds near-optimal solutions with much smaller ranks than the ones found cuLoRADS. For the largest instances considered in the table below the midrule line, cuHALLaR with HSLR format achieves 2x–4x speedups over cuLoRADS.
Problem Instance HALLaR vs cuHALLaR (on-the-fly) cuHALLaR vs cuLoRADS
Size $(n_{1},n_{2};m)$ $r$ CPU GPU Ratio cuHALLaR (HSLR) cuHALLaR (SDPA) cuLoRADS (SDPA) Ratio
3,000, 7,000; 828,931 3 7.63/3 0.77/3 9.9 3.33/3 3.22/3 123.71/19 37.1
3,000, 7,000; 828,931 3 7.52/3 0.70/3 10.7 3.20/3 3.25/3 11.42/19 3.6
3,000, 7,000; 2,302,586 5 34.52/5 1.28/5 27.0 3.37/5 2.49/5 3.19/19 0.9
3,000, 7,000; 2,302,586 5 34.42/5 1.09/5 31.6 2.45/5 2.68/5 2.26/19 0.9
10,000, 21,623; 2,948,996 3 38.88/3 1.22/3 31.9 3.31/3 2.68/3 5.14/21 1.5
10,000, 21,623; 2,948,996 3 34.80/3 1.11/3 31.4 4.10/3 2.94/3 5.65/21 1.4
10,000, 21,623; 8,191,654 5 141.17/5 2.05/5 68.9 4.02/5 7.38/5 4.71/21 1.2
10,000, 21,623; 8,191,654 5 140.55/5 1.95/5 72.1 4.02/5 7.74/5 9.00/21 2.2
25,000, 50,000; 3,367,574 2 43.52/2 1.20/2 36.3 4.40/2 3.96/2 10.77/23 2.4
25,000, 50,000; 7,577,040 3 79.93/3 1.82/3 43.9 4.19/3 3.68/3 21.31/23 5.1
30,000, 70,000; 4,605,171 2 60.57/2 1.63/2 37.2 4.43/2 3.48/2 17.02/24 3.8
30,000, 70,000; 10,361,633 3 145.76/3 2.57/3 56.7 5.12/3 4.54/3 18.45/24 3.6
50,000, 100,000; 7,151,035 2 121.63/2 1.89/2 64.4 12.36/2 5.18/2 23.67/24 1.9
50,000, 100,000; 16,089,828 3 242.00/3 3.68/3 65.8 7.04/3 5.69/3 27.95/24 4.0
80,000, 120,000; 9,764,859 2 237.38/2 2.45/2 96.9 5.63/2 6.05/2 20.66/25 3.7
80,000, 120,000; 21,970,931 3 490.07/3 5.02/3 97.6 6.92/3 6.27/3 32.28/25 4.7
120,000, 180,000; 34,051,152 3 1023.0/4 7.35/4 139.2 9.85/4 17.62/4 41.24/26 4.2
160,000, 240,000; 46,437,192 3 1419.26/4 8.79/4 161.5 12.24/3 12.12/3 40.65/26 3.3
240,000, 360,000; 71,845,299 3 1949.8/3 12.82/3 152.1 28.63/3 20.76/3 67.11/27 2.3
320,000, 480,000; 97,865,043 3 2754.83/3 17.41/3 158.2 36.76/3 31.78/3 136.43/28 3.7
400,000, 600,000; 124,339,596 3 3472.13/3 22.31/3 155.6 63.39/3 42.01/3 176.81/28 2.8
480,000, 720,000; 151,176,587 3 4403.04/3 28.13/3 156.5 71.02/3 46.21/3 159.7/28 2.2
560,000, 840,000; 178,314,984 3 6400.22/3 34.72/3 184.3 95.53/3 71.34/3 192.75/29 2.0
640,000, 960,000; 205,711,405 3 7374.16/3 39.25/3 187.9 96.83/3 100.77/3 319.17/29 3.3
720,000, 1,080,000; 233,333,416 3 8003.14/3 48.85/3 163.8 119.29/3 109.05/3 258.16/29 2.2
800,000, 1,200,000; 261,155,840 3 8771.27/3 53.3/3 164.6 114.11/3 110.11/3 293.65/30 2.6
@@
Table 2: Performance comparison between HALLaR, cuHALLaR, and cuLoRADS for matrix completion problems using a tolerance of $\epsilon=10^{-5}$. Time limit is set at three hours. Ratio is computed as cuLoRADS (SDPA) / cuHALLaR (HSLR).
@@ 5.3 Experiments for maximum stable set
Given a graph $G=([n],E)$, the maximum stable set problem consists of finding a subset of vertices of largest cardinality such that no two vertices are connected by an edge. Lovász introduced the $\vartheta$-function, which upper bounds the value of the maximum stable set. The $\vartheta$-function is the value of the following SDP relaxation
$\max_{X\in\mathbb{S}^{n}}\quad\{ee^{\top}\bullet X\;:\;X_{ij}=0,\;ij\in E,\;\mathrm{Tr}(X)\leq 1,\;X\succeq 0\}$ (33)
where $e=(1,1,\dots,1)^{\top}\in\mathbb{R}^{n}$ is the all ones vector. It was shown in [31] that the $\vartheta$-function agrees exactly with the stability number for perfect graphs.
@@
10,937; 75,488 wing_nodal DIMACS10 3,826.34/130 246.71/127 15.5
16,384; 49,122 delaunay_n14 DIMACS10 643.90/36 102.43/34 6.3
16,386; 49,152 fe-sphere DIMACS10 21.74/3 17.82/3 1.2
22,499; 43,858 cs4 DIMACS10 749.95/15 588.46/115 1.3
25,016; 62,063 hi2010 DIMACS10 2,607.24/15 132.85/21 19.6
@@
32,580; 77,799 vt2010 DIMACS10 2,952.84/18 196.59/60 15.0
48,837; 117,275 nh2010 DIMACS10 7,694.06/20 2,186.74/172 3.5
24,300; 34,992 aug3d GHS_indef 30.10/1 2.08/1 14.5
32,430; 54,397 ia-email-EU Network Repo 598.76/5 52.86/4 11.3
11,806; 32,730 Oregon-2 SNAP 1,581.11/23 467.75/55 3.4
21,363; 91,286 ca-CondMat SNAP 8,521.04/59 348.59/77 24.4
31,379; 65,910 as-caida_G_001 SNAP 2,125.41/8 428.41/27 5.0
26,518; 65,369 p2p-Gnutella24 SNAP 312.75/4 55.99/10 5.6
22,687; 54,705 p2p-Gnutella25 SNAP 242.42/4 88.53/13 2.7
36,682; 88,328 p2p-Gnutella30 SNAP 506.26/5 61.84/10 8.2
62,586; 147,892 p2p-Gnutella31 SNAP 1,484.71/5 245.33/29 6.0
49,152; 69,632 cca AG-Monien 63.09/2 12.30/2 5.1
49,152; 73,728 ccc AG-Monien 14.52/2 20.62/2 0.7
49,152; 98,304 bfly AG-Monien 15.72/2 10.35/2 1.5
16,384; 32,765 debr_G_12 AG-Monien 250.40/10 189.57/10 1.3
32,768; 65,533 debr_G_13 AG-Monien 471.80/9 638.54/13 0.7
65,536; 131,069 debr_G_14 AG-Monien 474.22/9 29.51/9 16.1
131,072; 262,141 debr_G_15 AG-Monien 521.88/10 28.07/10 18.6
262,144; 524,285 debr_G_16 AG-Monien 1,333.31/12 84.28/13 15.8
524,288; 1,048,573 debr_G_17 AG-Monien 6,437.05/12 171.81/13 37.4
1,048,576; 2,097,149 debr_G_18 AG-Monien 16,176.30/13 119.50/13 135.4
Table 3: Runtimes (in seconds) for the Maximum stable set problem. A relative tolerance of $\epsilon=10^{-5}$ is set. Ratio is computed as HALLaR/ cuHALLaR.
The results presented in Table 3 reveal a compelling scaling advantage of cuHALLaR over its CPU counterpart, HALLaR, particularly as the dimensions of the SDP instance increase. This trend is most pronounced within the AG-Monien dataset, where for de Bruijn graphs, the performance ratio grows significantly with the number of vertices. For instance, the speedup escalates from 16.1x for debr_G_14 to a remarkable 135.4x for debr_G_18, an instance with over one million vertices and over two million edges. For these large-scale problems, both solvers converge to solutions of nearly identical rank, indicating that the speedup is a direct consequence of the GPU’s superior computational throughput for cuHALLaR’s core linear algebra operations.
Performance patterns across other graph families confirm the robustness of this advantage. For the SNAP dataset, cuHALLaR delivers consistent and substantial improvements, with speedups ranging from 3.4x to 24.4x. Similarly, for the DIMACS10 instances, cuHALLaR achieves significant accelerations up to 19.6x compared to HALLaR.
Conversely, the results also show scenarios where GPU acceleration offers limited benefits. For smaller or computationally less demanding graphs, such as fe-sphere or ccc, the overhead associated with GPU kernel launches and memory management outweighs the parallelism benefits, leading to modest speedups or even slowdowns. Furthermore, on certain instances like ri2010 and nh2010, cuHALLaR converges to solutions of substantially higher rank than HALLaR e.g., 190 vs. 17 for ri2010. Even in these cases, the runtimes remain competitive, highlighting cuHALLaR’s capacity in efficiently managing high-rank factors.
Problem Instance HALLaR (on-the-fly) cuHALLaR (on-the-fly) cuLoRADS HALLaR/cuHALLaR cuLoRADS/cuHALLaR
@@
Table 4: Runtimes (in seconds) for the Maximum Stable Set problem on Hamming graphs. A relative tolerance of $\epsilon=10^{-5}$ is set. The HALLaR and cuHALLaR benchmarks were performed using on-the-fly problem generation.
Problem Instance On-the-fly Generation HSLR format
Graph($n$; $|E|$) HALLaR cuHALLaR Ratio cuHALLaR (HSLR)
$H_{10,2}$(1,024; 5,120) 0.153/2 1.735/2 0.088 0.798/2
$H_{11,2}$(2,048; 11,264) 0.260/2 1.627/2 0.160 0.92/2
$H_{12,2}$(4,096; 24,576) 0.606/2 2.033/2 0.298 1.83/2
$H_{13,2}$(8,192; 53,248) 1.852/2 3.106/2 0.596 1.10/2
$H_{14,2}$(16,384; 114,688) 3.843/2 3.170/2 1.212 1.28/2
$H_{15,2}$(32,768; 245,760) 8.240/2 3.223/2 2.557 2.701/2
$H_{16,2}$(65,536; 524,288) 17.235/2 4.457/2 3.869 3.83/2
$H_{17,2}$(131,072; 1,114,112) 33.342/2 3.904/2 8.541 2.51/2
$H_{18,2}$(262,144; 2,359,296) 46.436/2 2.627/2 17.681 3.43/2
$H_{19,2}$(524,288; 4,980,736) 105.408/2 5.281/2 19.958 10.70/2
$H_{20,2}$(1,048,576; 10,485,760) 249.701/2 8.109/2 30.786 6.62/2
$H_{21,2}$(2,097,152; 22,020,096) 592.696/2 16.966/2 34.934 12.96/2
$H_{22,2}$(4,194,304; 46,137,344) 1,702.255/2 37.145/2 45.826 18.842/2
$H_{23,2}$(8,388,608; 96,468,992) 4,205.741/2 83.905/2 50.125 51.75/2
Table 5: Runtimes (in seconds) for the Maximum Stable Set problem on Hamming graphs. A relative tolerance of $\epsilon=10^{-5}$ is set. Ratio in the fourth column is computed as HALLaR/ cuHALLaR. The last column provides two values in each line: the first one is the run time using the HSLR format and the second one is the rank of the solution.